# TikTok’s recommendations skewed towards Republican content during the 2024 U.S. presidential race

### v1

@@ Introduction
Considering this gap, we recompute the ideological skew measured in Figure 2G, once while only considering positive-partisan videos (Pro Democratic and Pro Republican), and again while only considering negative-partisan videos. For each of these sets of videos, we also recompute the expected ideological skew based on the different engagement metrics these videos receive (Supplementary Tables S16 and  S17 detail the ideological skews as measured by each of these models for positive-partisan and negative-partisan videos, respectively). The results of these analyses are presented in Figures 4C and  4D for positive- and negative-partisan videos, respectively. Broadly speaking, the Republican skew observed in positive-partisan videos is markedly smaller than that in negative-partisan videos. Formalizing this analysis using a logistic regression (Figure 4E) with video mismatch as an outcome variable suggests that negative-partisanship videos were 1.78 times more likely to be recommended as an ideological mismatch relative to positive-partisanship ones. See Supplementary Table S25 for the logistic regression table containing these results.
Topic analysis
What do partisan videos discuss, and what topics are most likely to be recommended as an ideological mismatch? To answer these questions, we use methods from [59] to identify the substantive topics covered in a given political video (see details in the Topic Analysis section of the Supplementary Materials). To start, we first isolate topics that were covered in at least 100 unique videos. We then compute the difference between the proportion of Republican- and Democratic-aligned videos on a given topic. Figure 5A illustrates differences in Republican and Democratic coverage of different topics, ordered from most proportionally Republican to most proportionally Democratic. As can be seen, topics stereotypically associated with the Democratic party (e.g., climate change, as well as abortion and reproductive health) emerge as topics with significantly more proportional coverage from Democratic-aligned videos. On the other hand, topics such as immigration, foreign policy, or the Ukraine war were topics more frequently covered by Republican-aligned videos. See Supplementary Table S20 for additional details regarding the distribution of topics within Democratic- and Republican-aligned videos, as well as chi-squared tests comparing these distributions.
@@
Having established that Democratic-conditioned bots viewed significantly more ideologically mismatched videos in previous sections, next, we analyze what topics such videos discussed. To this end, we focus on topics with a given partisan stance (Democratic- or Republican-aligned) that were viewed at least 100 times. For each such topic, Figure 5B illustrates the proportion of watches by Democratic-conditioned bots that were Democratic-aligned (blue) and Republican-aligned (red). Similarly, Figure 5C illustrates these proportions but for Republican-conditioned bots. Videos seen by Democratic-conditioned bots on the topics of immigration, crime, the Gaza conflict, or foreign policy broadly, were most likely to be of an opposing ideological alignment, with ideological mismatches representing the 56.9% and 58.9% of the latter two topics. In contrast, videos on the topics of abortion and reproductive health were proportionally the most commonly mismatched videos for Republican-conditioned bots, although these only account for 25.0% of all abortion-related videos seen by Republican bots. Supplementary Figures S8A and  S8B mirror Figures 5B and  5C while including videos classified as “Neutral” as well. Supplementary Figure S7 illustrates the difference in the proportion of videos on a topic with a given stance and the proportion of videos of that stance out of all political videos.
@@
Lastly, we also conduct an exploratory analysis of misinformation on TikTok. Using 20,950 “fake news” headlines collected from Snopes [60] and Politifact [61], we compute the cosine similarity of the vector embeddings of these headlines, and those of the video transcripts. As shown in Supplementary Table S9, at similarity thresholds above 0.7, less than 1% of videos of each ideological stance include misinformation as computed by this metric. See the Misinformation Classification section of the Supplementary material for more details.
@@ Discussion
Author Contributions
Competing Interests
The authors declare no competing interests.
Data and Code Availability
Supplementary Materials
Materials and Methods
@@
Over the duration of 27 weeks, 567 experiments were conducted. Specifically, each week, 21 new TikTok accounts are created by randomly combining the most common American first and last names from [70] and assigning an age between 22 and 24. This allowed each account to impersonate a potential voter for the U.S. presidential election likely to be active on TikTok. The 21 accounts created each week are split into one of nine experimental conditions, which are defined by two attributes. The first attribute is the state in which the bot is manually geo-located to, which is either New York, Texas, or Georgia [^1] Georgia was largely regarded as a key swing state for 2024 U.S. presidential elections, with the narrow 0.23% margin in favor of Joe Biden in the previous 2020 elections.. These states were chosen specifically as projected Democrat, Republican, and Swing states in the 2024 U.S. presidential elections, respectively. The second variable is the political leaning of the videos the bot watches in the “conditioning” stage. The videos watched in the conditioning stage are either published by known Democrat supporting channels or Republican supporting channels. Lastly, each week, three bots geo-located to Georgia bypass the conditioning stage of the experiment, and move directly to the recommendation stage. This is done to collect recommendations made to users who do not have a particular interest in politics. A summary of the experimental conditions can be seen in Supplementary Table S1 below and an illustration of the experimental setup can be seen in Figure 1 in the main text. Supplementary Figure S2 illustrates a more detailed timeline of a bot during a given experimental run.
@@
Table S1: Summary of experimental conditions
@@
In total, 54 candidate channels were compiled across both political party affiliations. Each channel associated with a given political party was then matched with a channel from the opposite political party based on its number of followers and cumulative number of likes using Euclidean distance matching. This is done to account for the possibility that the strength of a given bot’s political conditioning may be partially attributed to the popularity of the channels it watches in the conditioning stage. After matching, the 10 pairs of channels with the smallest Euclidean distance were selected. Supplementary Table S3 below details the channels used in the conditioning stage, showing 12 pairs in total, as two channels (“donaldtrumpwasright”, “kayetriots”) became no longer available (deleted or made private) during our experiments. As such, those two pairs were replaced with the new pairs that had the smallest Euclidean distance. The accounts of the main political candidates (kamalaharris and realdonaldtrump) did not exist when the experiment began in May, and hence, were not included as conditioning channels. Kamala Harris officially joined TikTok in July, while Donald Trump joined the platform in June.
In each experiment, a bot is conditioned to lean either Democratic or Republican by watching up to 50 most recent videos from 8 randomly selected channels aligned with the respective political party. The target of each bot watching 400 conditioning videos in total was not always achieved, as shown in Supplementary Table S2. On average, Republican bots viewed less conditioning videos due to several of the Republican accounts having less than 50 total published videos, particularly during the first few months of the experiment. To account for this, the number of videos watched during the conditioning stage was controlled for during all analyses that compared recommendation rates across experimental conditions. Every video during the conditioning stage is watched for one minute, ensuring consistent exposure to the content from the selected channels. After completing the predefined set of videos, the bot would then “sleep” for 24 hours, where it would not open or interact with TikTok. This pause was implemented to simulate realistic human viewing patterns, minimizing the risk of bot detection that could result from consuming an excessive number of videos in rapid succession.
@@
Table S2: The mean number of conditioning videos watched per experimental condition.
@@
Table S3: Left- and right-leaning TikTok accounts used during the conditioning phase, paired based on the minimum Euclidean distance with respect to the number of followers and likes.
@@
There were a handful of weeks in which the bots failed to meet our inclusion threshold. Firstly, during the end of July and the month of August, bots with a Texas geo-location lost internet connectivity while the authors were not available to tend to the bots. As such, we were unable to collect data from Texas during this time. In another weeks, some accounts were recognized as bots by TikTok’s bot detection algorithm and subsequently suspended before reaching 150 video threshold. Of the 567 experiments conducted, 323 met our inclusion criteria. Supplementary Table S11 below lists the total number of bots which met the inclusion criteria on a weekly basis across the different bot conditioning and geo-location classes.
@@
The distribution of classifications across these categories is presented in Supplementary Table S4 below.
@@
Table S4: The number of videos classified as political, pertaining to the U.S. presidential election, and of each ideological stance.
@@
To ensure classification reliability, we employed a consensus-based approach where GPT-4 served as the tiebreaker in cases of disagreement between GPT-4o and Gemini-Pro. The inter-model agreement rates for each classification task are detailed in Supplementary Table S4 below. This ensemble method was chosen to mitigate individual model biases and enhance the robustness of our ideological stance classifications.
@@
Table S5: Inter-LLM agreement metrics for the three prompts.
@@
To validate our automated classification approach, we conducted a human annotation study on a subset of 500 randomly selected transcripts. Three independent annotators, all political science undergraduate students, were recruited to perform the classification tasks. The annotators were permitted to use web searches to research unfamiliar topics or references within the transcripts, ensuring informed labeling decisions. Each transcript received independent classifications from all three annotators following the same categorical framework used in the LLM classification. The results of this validation study, including inter-rater reliability metrics, human majority-LLM majority classification accuracy, Cohen’s kappa coefficient comparing human majority and LLM majority decisions, and the LLM ensemble’s F1-scores, are presented in Supplementary Table S6, while Supplementary Table S7 details these scores for each model separately.
@@
Table S6: Human validation accuracy and agreement metrics for LLM-ensemble classification tasks.
@@
Table S7: The accuracy, Cohen kappa, and F1-scores of each LLM model individually, as well as the LLM majority vote, when compared with human ground truth labels.
@@
To identify channels potentially leaning toward either the Democratic or Republican ideologies, we focused on channels that had published videos previously labeled with a specific non-neutral ideological stance (e.g., “Anti-Democrat,” “Pro-Republican,” “Anti-Republican,” or “Pro-Democrat”). From this process, we identified 170 unique channels for analysis. For each channel, we calculated the proportion of videos aligned with a particular party’s ideology as a fraction of the total videos collected for that channel. To ensure robustness in our classification, channels were labeled as Democrat-aligned or Republican-aligned if more than 75% of their analyzed videos were ideologically consistent with one party. For 85 channels with fewer than 10 unique videos collected during the experiment phase, we used TikAPI [57] to retrieve up to 30 additional videos published prior to the election. These additional videos were processed through the same transcript classification pipeline to determine their ideological stance. This step was necessary to avoid potential misclassification caused by a limited number of channels’ unique videos watched by the bots. Table S8 presents a summary of the classified channels, including the total number of channels in each category (Republican-aligned, Democrat-aligned, and neutral) as well as the average proportion of partisan content for each group.
@@
Table S8: Classification of TikTok channels based on political videos, including channel counts and average proportions of partisan content.
@@
Given that our primary analysis throughout this study focuses on videos with transcripts, we must ensure that this sample is representative of videos throughout the entire dataset. To do so, we annotate a random sample of videos recommended to each of the set of Republican and Democrat during the experiment. Specifically, for each of these sets, we analyze a random sample of 2000 videos which do not include transcripts to measure the following features of a given video: (1) whether the video is political in nature, (2) whether the video is concerned with the 2024 U.S. elections or major U.S. political figures, and (3) the ideological stance of the video in question (Pro-Democrat, Anti-Democrat, Pro-Republican, Anti-Republican, or Neutral). This annotation task was again completed by three political science undergraduates, with high agreement between annotators (Krippendorf’s alpha; Q1: $\alpha=0.858$, Q2: $\alpha=1.0$, Q3: $\alpha=0.989$). Indeed, we find that the inter-rater agreement when viewing the videos directly, rather than relying on the transcript alone (the results of which are in Supplementary Table S6), yielded higher agreement between the raters.
This process allows us to compare the distribution of political content in videos with transcripts against those without transcripts. The results of this comparison is illustrated in Supplementary Figure S1. As can be seen, videos with no transcripts contained significantly fewer political videos on average relative to those that did include transcripts. This result is perhaps unsurprising given that videos which do not include a transcript are also less likely to include dialogue (e.g., videos of landscapes, pets, etc.). However, of those which were political, we do not find significant differences (computed through chi-squared tests) between transcript and non-transcript videos with regards to the proportion of political videos pertaining to the election, nor their distribution of Democrat-aligned, Republican-aligned, or Neutral videos.
Figure S1: (A, B) A comparison between videos with and without transcripts seen by Democrat and Republican bots, respectively.
@@
Using these vector embeddings, we compute cosine similarities across all possible headline-transcript pairs. Supplementary Table S9 lists the percentage of transcripts of a given ideological stance which had a headline with a cosine similarity greater than a certain threshold. As can be seen, at thresholds greater than 0.8, less than 1% of Anti Democrat and Neutral videos include misinformation, and no Anti Republican, Pro Democrat, or Pro Republican videos included misinformation as computed by this metric. While this process mirrors approaches taken by prior work analyzing misinformation on social media [29, 75], further efforts that rely on larger datasets of misinformation, especially those related to the 2024 U.S. elections, may be needed to fully capture the extent of misinformation on TikTok.
@@
Table S9: The percentage of transcripts that had a headline cosine similarity score above a certain threshold.

### v2

@@ Introduction
Having established that Democratic-conditioned bots viewed significantly more ideologically mismatched videos in previous sections, next, we analyze what topics such videos discussed. To this end, we focus on topics with a given partisan stance (Democratic- or Republican-aligned) that were viewed at least 100 times. For each such topic, Figure 6B illustrates the proportion of watches by Democratic-conditioned bots that were Democratic-aligned (blue) and Republican-aligned (red). Similarly, Figure 6C illustrates these proportions but for Republican-conditioned bots. Videos seen by Democratic-conditioned bots on the topics of immigration, crime, the Gaza conflict, or foreign policy broadly, were most likely to be of an opposing ideological alignment, with ideological mismatches representing the 56.9% and 58.9% of the latter two topics. In contrast, videos on the topics of abortion and reproductive health were proportionally the most commonly mismatched videos for Republican-conditioned bots, although these only account for 25.0% of all abortion-related videos seen by Republican bots. Supplementary Figures 7A and 7B mirror Figures 6B and  6C while including videos classified as “Neutral” as well. Supplementary Figure 6 illustrates the difference in the proportion of videos on a topic with a given stance and the proportion of videos of that stance out of all political videos.
@@
Lastly, we also conduct an exploratory analysis of misinformation on TikTok. Using 20,950 “fake news” headlines collected from Snopes [60] and Politifact [61], we compute the cosine similarity of the vector embeddings of these headlines, and those of the video transcripts. As shown in Table 12, at similarity thresholds above 0.7, less than 1% of videos of each ideological stance include misinformation as computed by this metric. See the Misinformation Classification section of the Methods for more details.
@@ Discussion
For the survey component of this study, the survey was deemed exempt by the authors’ institutional review board (IRB Protocol Number: HRPP-2025-69)
Methods
@@
Over the duration of 27 weeks, 567 experiments were conducted. Specifically, each week, 21 new TikTok accounts are created by randomly combining the most common American first and last names from [70] and assigning an age between 22 and 24. This allowed each account to impersonate a potential voter for the U.S. presidential election likely to be active on TikTok. The 21 accounts created each week are split into one of nine experimental conditions, which are defined by two attributes. The first attribute is the state in which the bot is manually geo-located to, which is either New York, Texas, or Georgia [^1] Georgia was largely regarded as a key swing state for 2024 U.S. presidential elections, with the narrow 0.23% margin in favor of Joe Biden in the previous 2020 elections.. These states were chosen specifically as projected Democrat, Republican, and Swing states in the 2024 U.S. presidential elections, respectively. The second variable is the political leaning of the videos the bot watches in the “conditioning” stage. The videos watched in the conditioning stage are either published by known Democrat supporting channels or Republican supporting channels. Lastly, each week, three bots geo-located to Georgia bypass the conditioning stage of the experiment, and move directly to the recommendation stage. This is done to collect recommendations made to users who do not have a particular interest in politics. A summary of the experimental conditions can be seen in Table 2. Supplementary Figure 1 illustrates a more detailed timeline of a bot during a given experimental run.
@@
Table 2: Summary of experimental conditions
@@
In total, 54 candidate channels were compiled across both political party affiliations. Each channel associated with a given political party was then matched with a channel from the opposite political party based on its number of followers and cumulative number of likes using Euclidean distance matching. This is done to account for the possibility that the strength of a given bot’s political conditioning may be partially attributed to the popularity of the channels it watches in the conditioning stage. After matching, the 10 pairs of channels with the smallest Euclidean distance were selected. Table 4 below details the channels used in the conditioning stage, showing 12 pairs in total, as two channels (“donaldtrumpwasright”, “kayetriots”) became no longer available (deleted or made private) during our experiments. As such, those two pairs were replaced with the new pairs that had the smallest Euclidean distance. The accounts of the main political candidates (kamalaharris and realdonaldtrump) did not exist when the experiment began in May, and hence, were not included as conditioning channels. Kamala Harris officially joined TikTok in July, while Donald Trump joined the platform in June.
In each experiment, a bot is conditioned to lean either Democratic or Republican by watching up to 50 most recent videos from 8 randomly selected channels aligned with the respective political party. The target of each bot watching 400 conditioning videos in total was not always achieved, as shown in Table 3. On average, Republican bots viewed less conditioning videos due to several of the Republican accounts having less than 50 total published videos, particularly during the first few months of the experiment. To account for this, the number of videos watched during the conditioning stage was controlled for during all analyses that compared recommendation rates across experimental conditions. Every video during the conditioning stage is watched for one minute, ensuring consistent exposure to the content from the selected channels. After completing the predefined set of videos, the bot would then “sleep” for 24 hours, where it would not open or interact with TikTok. This pause was implemented to simulate realistic human viewing patterns, minimizing the risk of bot detection that could result from consuming an excessive number of videos in rapid succession.
@@
Table 3: The mean number of conditioning videos watched per experimental condition.
@@
Table 4: Left- and right-leaning TikTok accounts used during the conditioning phase, paired based on the minimum Euclidean distance with respect to the number of followers and likes.
@@
There were a handful of weeks in which the bots failed to meet our inclusion threshold. Firstly, during the end of July and the month of August, bots with a Texas geo-location lost internet connectivity while the authors were not available to tend to the bots. As such, we were unable to collect data from Texas during this time. In another weeks, some accounts were recognized as bots by TikTok’s bot detection algorithm and subsequently suspended before reaching 150 video threshold. Of the 567 experiments conducted, 323 met our inclusion criteria. Supplementary Table 1 lists the total number of bots which met the inclusion criteria on a weekly basis across the different bot conditioning and geo-location classes.
@@
The distribution of classifications across these categories is presented in Table 5 below.
@@
Table 5: The number of videos classified as political, pertaining to the U.S. presidential election, and of each ideological stance.
@@
To ensure classification reliability, we employed a consensus-based approach where GPT-4 served as the tiebreaker in cases of disagreement between GPT-4o and Gemini-Pro. The inter-model agreement rates for each classification task are detailed in Table 6 below. This ensemble method was chosen to mitigate individual model biases and enhance the robustness of our ideological stance classifications.
@@
Table 6: Inter-LLM agreement metrics for the three prompts.
@@
To validate our automated classification approach, we conducted a human annotation study on a subset of 500 randomly selected transcripts. Three independent annotators, all political science undergraduate students, were recruited to perform the classification tasks. The annotators were permitted to use web searches to research unfamiliar topics or references within the transcripts, ensuring informed labeling decisions. Each transcript received independent classifications from all three annotators following the same categorical framework used in the LLM classification. The results of this validation study, including inter-rater reliability metrics, human majority-LLM majority classification accuracy, Cohen’s kappa coefficient comparing human majority and LLM majority decisions, and the LLM ensemble’s F1-scores, are presented in Table 7, while Table 8 details these scores for each model separately.
@@
Table 7: Human validation accuracy and agreement metrics for LLM-ensemble classification tasks.
@@
Table 8: The accuracy, Cohen kappa, and F1-scores of each LLM model individually, as well as the LLM majority vote, when compared with human ground truth labels.
@@
To identify channels potentially leaning toward either the Democratic or Republican ideologies, we focused on channels that had published videos previously labeled with a specific non-neutral ideological stance (e.g., “Anti-Democrat,” “Pro-Republican,” “Anti-Republican,” or “Pro-Democrat”). From this process, we identified 170 unique channels for analysis. For each channel, we calculated the proportion of videos aligned with a particular party’s ideology as a fraction of the total videos collected for that channel. To ensure robustness in our classification, channels were labeled as Democrat-aligned or Republican-aligned if more than 75% of their analyzed videos were ideologically consistent with one party. For 85 channels with fewer than 10 unique videos collected during the experiment phase, we used TikAPI [57] to retrieve up to 30 additional videos published prior to the election. These additional videos were processed through the same transcript classification pipeline to determine their ideological stance. This step was necessary to avoid potential misclassification caused by a limited number of channels’ unique videos watched by the bots. Table 9 presents a summary of the classified channels, including the total number of channels in each category (Republican-aligned, Democrat-aligned, and neutral) as well as the average proportion of partisan content for each group.
@@
Table 9: Classification of TikTok channels based on political videos, including channel counts and average proportions of partisan content.
Comment Classification
To identify the ideological stance of the comments on a given TikTok video, we follow a similar methodology to that of classifying a video. Specifically, we pass both the transcript of the video, as well as the comment in question in a prompt to GPT-4o, which can be seen below:
”Given the following TikTok video transcript and comment, classify the comment into one of the following categories.
Transcript of the video: TRANSCRIPT
Comment: COMMENT
Categories:
Anti Democrat
Anti Republican
Pro Democrat
Pro Republican
Neutral”
The outputs of the model were validated by three independent annotators, again all political science undergraduate students. The results of this validation process are shown in Table 10 below.
Inter-Rater Reliability
Krippendorf’s Alpha
Accuracy Cohen Kappa F1 Score
Comment Classification 0.89371 0.95016 0.93366 0.95025
Table 10: Human validation accuracy and agreement metrics for comment classification task.
Sensitivity Analysis
To effectively test the strength of the bias observed in our analysis, we conduct a sensitivity analysis with regards to the various engagement metrics analyzed in our study. Specifically, here, we ask the following question. If the observed bias was due to some unobserved internal engagement metric only available to TikTok, how different would Republican and Democrat videos need to be in this engagement metric to explain the observed biases measured in our study. To answer this question, we formulate this unobserved metric as a random variable sampled from a set of different distributions, namely, a binomial, lognormal, normal, or poisson distribution. Democratic videos and Republican videos then sample from two distributions of a given type which differ based only on their means. As an example, Democratic videos would sample from a normal distribution with a mean of 1, while Republican videos would sample from a normal distribution with a mean of 2. We can then ask, how large of a difference in means between these two distributions would result in a bias consistent with that seen in our study. The results of this analysis can be seen in Table 1 below.
Specifically, in Table 11 below, we show in the first column the true difference between the Republican and Democratic videos with regards to their normalized values for a given engagement metric. In the remaining four columns, we show the ratio between the difference needed to match the observed bias and that of the true difference in mean between Republican and Democratic videos in a given engagement metric. As can be seen, for the majority of engagement metrics, it is Democratic videos which tend to have higher engagement on average, with the exceptions being in the video comment count, like count, play count, share count, and composite score. Although the combination of these metrics produces the largest normalized Republican skew in engagement gap, this gap would need to be 9.4 times as large in order to full explain the observed bias in recommendation rates seen in our experiments.
Required Scaling with different distributions
Engagement metric
True diff.
in normalized
metric values
(Rep - Dem)
Binomial Lognormal Normal Poisson
(Comments, Likes, Shares, Plays) Score 0.0209 31.2593 9.4323 34.0259 29.03865
Video share count 0.0028 169.4288 51.1241 184.4241 157.3927
Video play count 0.0024 177.4398 53.5413 193.1442 164.8346
Comment count 0.0006 782.0713 235.9851 851.2887 726.5135
Video Like count 0.0009 325.8119 98.3117 354.6479 302.6664
Comments per recommendation -0.0005 -77.8383 -23.4872 -84.7274 -72.3087
Video engagement rate -0.0016 -70.573 -21.295 -76.8191 -65.5596
Channel follower count -0.0017 -122.0704 -36.834 -132.8742 -113.3986
Likes per recommendation -0.0026 -73.7079 -22.2409 -80.2314 -68.4717
Channel cumul. likes -0.0116 -48.4503 -14.6196 -52.7384 -45.0084
Video length -0.014 -31.1608 -9.4026 -33.9187 -28.9472
Channel video count -0.0143 -26.1214 -7.882 -28.4333 -24.2657
Opposing comment proportion -0.0543 -15.3969 -4.6459 -16.7596 -14.3031
Party-aligned comment proportion -0.1043 -4.8802 -1.4726 -5.3121 -4.5335
Verified account -0.1556 -3.6068 -1.0883 -3.926 -3.3505
Table 11: The true difference between Republican and Democratic videos with regards to their normalized values for a given engagement metric as well as the ratio between the difference needed to match the observed bias and that of the true difference in mean between Republican and Democratic videos for a given engagement metric.
@@
Given that our primary analysis throughout this study focuses on videos with transcripts, we must ensure that this sample is representative of videos throughout the entire dataset. To do so, we annotate a random sample of videos recommended to each of the set of Republican and Democrat during the experiment. Specifically, for each of these sets, we analyze a random sample of 2000 videos which do not include transcripts to measure the following features of a given video: (1) whether the video is political in nature, (2) whether the video is concerned with the 2024 U.S. elections or major U.S. political figures, and (3) the ideological stance of the video in question (Pro-Democrat, Anti-Democrat, Pro-Republican, Anti-Republican, or Neutral). This annotation task was again completed by three political science undergraduates, with high agreement between annotators (Krippendorf’s alpha; Q1: $\alpha=0.858$, Q2: $\alpha=1.0$, Q3: $\alpha=0.989$). Indeed, we find that the inter-rater agreement when viewing the videos directly, rather than relying on the transcript alone (the results of which are in Table 7), yielded higher agreement between the raters.
This process allows us to compare the distribution of political content in videos with transcripts against those without transcripts. The results of this comparison is illustrated in Figure 7. As can be seen, videos with no transcripts contained significantly fewer political videos on average relative to those that did include transcripts. This result is perhaps unsurprising given that videos which do not include a transcript are also less likely to include dialogue (e.g., videos of landscapes, pets, etc.). However, of those which were political, we do not find significant differences (computed through chi-squared tests) between transcript and non-transcript videos with regards to the proportion of political videos pertaining to the election, nor their distribution of Democrat-aligned, Republican-aligned, or Neutral videos.
Figure 7: (A, B) A comparison between videos with and without transcripts seen by Democrat and Republican bots, respectively.
@@
Using these vector embeddings, we compute cosine similarities across all possible headline-transcript pairs. Table 12 lists the percentage of transcripts of a given ideological stance which had a headline with a cosine similarity greater than a certain threshold. As can be seen, at thresholds greater than 0.8, less than 1% of Anti Democrat and Neutral videos include misinformation, and no Anti Republican, Pro Democrat, or Pro Republican videos included misinformation as computed by this metric. While this process mirrors approaches taken by prior work analyzing misinformation on social media [29, 75], further efforts that rely on larger datasets of misinformation, especially those related to the 2024 U.S. elections, may be needed to fully capture the extent of misinformation on TikTok.
@@
Table 12: The percentage of transcripts that had a headline cosine similarity score above a certain threshold.